# Microscopic Analysis on LLM players via Social Deduction Game

### v1

@@ IV-C2 Result
Tables II (c) and III (c) elaborate measurements of camouflage when playing against strong and weak citizens, respectively. We excluded AR from the tables because there was no accusation during gameplay, regardless of the citizen type. Similarly, some values are missing for VR and VE because the spy ended the game early without votes by guessing the hidden location.
The results indicate that the variants of GPT have sufficient camouflage ability compared to other models. Against strong citizens, GPT-3.5 was the best model in terms of CR, followed by Gemini, GPT-4, and LlaMA2 (7.1, 22.2, 25.0, and 100%, respectively). The results for VR showed a similar tendency: GPT-3.5 (25%), GPT-4 (33.3%), Gemini (35.7%), and LlaMA2 (45.2%). Finally, regarding VE, GPT-4 was the best model, followed by GPT-3.5, LlaMA2, and Gemini (0.953, 0.730, 0.634, and 0.348, respectively). Interestingly, GPT-3.5 showed camouflage performance similar to that of GPT-4, although GPT-4 showed much higher performance in other language-based tasks[4].
Against weak citizens, the GPT variants showed good performance compared to others. Because the GPT variants ended the game by guessing, citizens had no chance to find the spy. Thus, the CR values for GPT-3.5 and GPT-4 were zero, and VR and VE could not be measured using these models. Other models showed a similar tendency to that against strong citizens. Regarding CR and VR, Gemini (0% and 7.1%, respectively) was better than LLaMA2 (60% and 13.5%, respectively). However, regarding VE, Gemini (0.420) performed worse than LLaMA2 (0.605), as shown by the results against strong citizens.
@@ IV-C3 Discussion
In this section, we discuss how to derive insights using the proposed metrics to conduct a multifaceted camouflage analysis. This discussion includes the following two aspects: (1) a complementary relationship between Vote Rate (VR) and Vote Entropy (VE) and (2) the need for formative camouflage metrics.
First, VR and VE are complementary metrics that enable a multifaceted analysis of camouflage. As suggested above, these two metrics can form three partitions explaining the different levels of camouflage ability. Using these partitions, we can determine the best model for balancing the two sub-abilities of camouflage; the GPT family showed gameplay balanced between VR and VE. These models can evade and disperse suspicion to win the game. In contrast, the worst model in terms of camouflage was LLaMA2. Although LLaMA2 performed better against weak citizens than strong citizens, LLaMA2 player was often suspected as a spy and disregarded the fact that the majority of players doubted him/her. Here, the two sub-abilities are both good or both bad in these two cases. Gemini showed a different behavior: the model was good at evading suspicion but struggled to disperse suspicion. Thus, as illustrated above, VR and VE can provide a wider picture of camouflage behavior.
Second, the proposed metrics provide additional perspectives for investigating camouflage compared to simple summative metrics. In Section IV-A, we observed a discrepancy between two summative metrics, winning rate (WR) and living round (LR). For example, both GPT-3.5 and LLaMA2 showed relatively low WR, but the performance gap regarding LR between them is large. Therefore, using only summative metrics, we can conclude that LLaMA2 outperformed GPT-3.5 in SpyGame. However, our metrics suggested different results; GPT-3.5 is good at camouflage, but LLaMA2 is not. This example shows that the proposed metrics are more suitable for examining camouflage ability than existing summative metrics.
Second, proposed metrics provide additional perspectives on investigating Camouflage, compared to simple summative metrics. In section IV-A, we observed a discrepancy between two summative metrics, winning rate (WR) and living round (LR). For example, both GPT-3.5 and LLaMA2 showed relatively low WR scores, but regarding LR, achieved relatively low and high performance, respectively. So, using summative metrics only, we may conclude that LLaMA2 outperforms GPT-3.5 in SpyGame. However, the proposed metrics suggested different results; GPT-3.5 is good at Camouflage, but LLaMA2 is not. This example shows that proposed metrics are more suitable to examine Camouflage ability than existing summative metrics.
Similarly, camouflage metrics are helpful in identifying gameplay problems. For example, both GPT-4 and GPT-3.5 showed relatively low LR scores; however, WR achieved the best and worst scores, respectively. Therefore, using only summative metrics, we can conclude that GPT-4 outperforms GPT-3.5 in SpyGame. However, the proposed metrics suggest different results: both GPT-4 and GPT-3.5 are good at camouflage regarding VR and VE.
@@ V Qualitative Analysis
To reconfirm the findings of the quantitative analysis from another perspective, we conducted a thematic analysis. We defined three types of reasoning errors in LLM reasoning while deciding their actions by coding abnormal patterns of 15 randomly selected gameplay logs. As a result, we found that these patterns are related to psychological concepts, including identity confusion, memory distortion, and dissociation[23].
In addition, we found that these concepts are related to the results of the quantitative analysis, especially for caught rate (CR) and guess success (GS). The following subsections illustrate the procedure and results of this qualitative analysis.
@@ V-A Analysis Procedure
$N$ Role ambiguity Memory distortion Dissociation
@@
The total number of reasoning steps in each condition
@@ V-B1 Exposure
Exposure is a case in which a citizen’s utterance literally reveals the hidden location. The Fleiss’ Kappa value for this category was 0.7875. Note that this pattern can only be identified in utterances rather than in reasoning steps. Therefore, we report the frequency of this category based on utterances; among 729 citizen utterances, 125 utterances were labeled as exposure. The following quote provides a representative example:
@@
We suspect that this exposure pattern is because LLMs may not have sufficient common sense to understand the game rules. In particular, SpyGame has a rule “citizens should not expose the location to the spy.” A human citizen player can notice that a direct mention to other players of the hidden location during a game causes their defeat. However, some LLMs may not have this common sense among players; they simply go through the mentions. Therefore, LLMs believe that they can reveal their information to other citizens because the game rule does not prohibit such behaviors and the citizens already know that fact. Thus, citizen LLMs sometimes publicly reveal their hidden information.
V-B2 Role ambiguity
Role ambiguity[25] is a case in which the spy confuses their own identity with that of a citizen. In an SDG, understanding a player’s identity is related to two major concepts affecting gameplay: (1) the team of the player and (2) the goal of the player. Therefore, we divided abnormal patterns of role ambiguity into two sub-categories: team misunderstanding and goal misunderstanding. These two subcategories can occur independently. We present details of the subcategories in the below paragraphs. In summary, Fleiss’ Kappa for this category was 0.563. Among 506 reasoning steps in the data, we identified 21 steps with role ambiguity.
@@ V-B2 Role ambiguity
We suspect that this category occurs because of the memory limitation of LLMs. To understand a player’s identity correctly, an LLM should retain information about the player’s identity during the decoding process. Since we provided the game rule, identity of a player, and conversation logs together as an input prompt, the LLM needs to distinguish information about the identity from the input prompt. However, such a distinction is difficult for the LLM because the other parts of the input prompt are much larger than the identity part. For example, the explanation of the game rules takes more than 130 words, whereas that of identity takes only 10 words. Thus, LLM players can confuse their identities with other characters written in the game rules, triggering role ambiguity.
V-B3 Memory distortion
V-B4 Dissociation
[Player4, the spy, tries to guess the hidden location.]
We suspect that this category occurs because LLMs are not good at differentiating between numerical symbols. As previous studies have reported that LLMs suffer from biases induced by proper names[27], we named each player as ‘Player#’ (e.g., Player1) to avoid such biases. Therefore, when an LLM attempts to collect clues from the conversation log, it has to distinguish the source of each utterance using a single difference between player names: a single digit. However, previous studies have reported that LLMs struggle to recognize different digit tokens during their computation processes. Thus, LLM players can misinterpret their words as those of others in the conversation log, which triggers dissociation.
@@ V-C Reconfirmation of Quantitative Analysis
Combining the results of the two analyses confirms the findings of the quantitative analysis. Specifically, we discovered that these quantitative labels, except exposure, can be connected to two abilities in the quantitative analysis: camouflage and intent identification.
First, role ambiguity increases the likelihood of a spy’s camouflage failure. When we compared the frequency of role ambiguity and the results of the three camouflage metrics, the trend of frequency was similar to that of the metrics. For example, against strong citizens, GPT-3.5 showed the lowest errors in role ambiguity and achieved the highest performance in camouflage metrics. Similarly, LLaMA2 showed more errors while achieving a lower performance than the other models. We suspect that this similarity between the trends is due to the loss of the spy’s need for camouflage. When role ambiguity occurs in the spy’s reasoning steps, the spy confuses their role with that of a citizen. Such confusion leads the spy to make an unusual utterance, which draws suspicion.
Second, memory distortion and dissociation may hinder a spy’s intent identification during the game. When we compared the frequency of memory distortion and dissociation with the results of the four intent identification metrics, the trend of frequency was similar to that of the metrics. For example, against strong citizens, GPT-3.5 showed the highest errors in both memory distortion and dissociation and achieved the lowest performance in intent identification. Similarly, GPT-4 showed fewer errors while achieving a higher performance than the other models. We suspect that this similarity between trends is due to the quality of information that the spy considers during their reasoning process. The spy player should select and combine pieces of information from the conversation to identify the hidden location. When memory distortion occurs, the spy can regard misinterpreted pieces as appropriate because they distorted the information. Similarly, when dissociation occurs, the spy regards the apocryphal information that they previously said as information provided by citizens. Therefore, such misunderstood pieces can be obstacles to reaching the correct answer.
@@ VI Conclusion
To assess whether LLMs can effectively support tricky conversations used in SDGs, we tackled two key issues with previous evaluation methodologies. The first is the reliance on macroscopic quantitative evaluations, which overlook the need for microscopic formative assessment within quantitative analyses. The second is the lack of non-systematic qualitative evaluation, highlighting the necessity for thematic analysis to uncover the logical reasoning behind gameplay. In our demonstration, we chose SpyGame, a variant of SpyFall, as the experimental environment. While human gameplay often incorporates nonverbal and paraverbal cues to identify the hidden intention of others, we adopted the simplest implementation of LLM players to simplify our demonstration. We analyzed the gameplay behaviors using four different LLMs in SpyGame logs. Our analysis was conducted in two ways. First, through quantitative analysis, we employed eight metrics to address the macroscopic evaluation issue. This analysis revealed that our metrics are more effective than existing macroscopic metrics for explaining intent identification and camouflage behavior. Second, in qualitative analysis, we performed thematic analysis to resolve the non-systematic evaluation issue. From the analysis, we discovered four major categories that affect LLMs’ gameplay as a spy. These major categories supported the findings of our quantitative analysis. Despite the limitations of our study, our work underscores the significance of microscopic analysis in evaluating LLMs’ gameplay. In the future, we plan to extend this work to multi-modal LLMs to better capture the differences between human and LLM players. We hope that this study provides valuable insights into the behavior of LLM-based players in SDGs.
VII Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2021-0-01341, Artificial Intelligence Graduate School Program, Chung-Ang University) and the Chung-Ang University Graduate Research Scholarship in 2024.

### v2

@@ IV-C2 Result
Tables II (c) and III (c) elaborate measurements of deceptive control when playing against strong and weak citizens, respectively. Note that some missing values for VR and VE indicate that the spy guessed the hidden location for all games, and no votes were conducted.
The results indicate that the variants of GPT have sufficient ability compared to other models. Against strong citizens, GPT-3.5 was the best model in terms of CR, followed by Gemini, GPT-4, and Llama (7.1, 22.2, 25.0, and 100%, respectively). The results for VR showed a similar tendency: GPT-3.5 (25%), GPT-4 (33.3%), Gemini (35.7%), and Llama (45.2%). Finally, regarding VE, GPT-4 was the best model, followed by GPT-3.5, Llama, and Gemini (0.953, 0.730, 0.634, and 0.348, respectively). Interestingly, GPT-3.5 showed deceptive control performance similar to that of GPT-4, although GPT-4 showed much higher performance in other language-based tasks[1].
Against weak citizens, GPT variants showed good performance compared to others. Because GPT variants ended the game by guessing, citizens had no chance to find the spy. Thus, CR values for GPT-3.5 and GPT-4 were zero, and VR and VE could not be measured using these models. Other models showed a tendency similar to that against strong citizens. Regarding CR and VR, Gemini (0% and 7.1%, respectively) was better than Llama (60% and 13.5%). However, regarding VE, Gemini (0.420) performed worse than Llama (0.605), as shown by results against strong citizens.
@@ IV-C3 Discussion
In this section, we discuss how to derive insights using the proposed metrics to conduct a multifaceted analysis of deceptive control. This discussion includes the following two aspects: (1) a complementary relationship between Vote Rate (VR) and Vote Entropy (VE), and (2) additional perspectives captured by adopting formative metrics.
First, VR and VE are complementary metrics that enable a multifaceted analysis of deceptive control. As suggested above, these two metrics can form three partitions explaining the different levels of capacity regarding deceptive control. Using these partitions, we can determine the best model for balancing two sub-abilities of deceptive control; GPT family showed gameplay balanced between VR and VE. These models can evade and disperse suspicion to win the game. In contrast, the worst model in terms of deceptive control was Llama. Although Llama performed better against weak citizens than strong citizens, Llama player was often suspected as a spy and disregarded the fact that the majority of players doubted him/her. So, the two sub-abilities are both good or both bad in these two cases. Gemini showed a different behavior: the model was good at evading suspicion but struggled to disperse suspicion. Thus, as illustrated above, VR and VE can provide a wider picture of deceptive control behavior.
Second, proposed metrics provide additional perspectives on investigating deceptive control, compared to simple summative metrics. In section IV-A, we observed a discrepancy between two summative metrics, win rate (WR) and living round (LR). For example, both GPT-3.5 and Llama showed relatively low WR scores, but regarding LR, they achieved relatively low and high performance, respectively. So, using summative metrics only, we may conclude that Llama outperforms GPT-3.5 in SpyGame. However, the proposed metrics suggested different results; GPT-3.5 is good at controlling information, but Llama is not. This example shows that our proposed metrics are more suitable for examining deceptive control ability than existing summative metrics.
Similarly, our proposed metrics are helpful in identifying gameplay problems. For example, both GPT-4 and GPT-3.5 showed relatively low LR scores; however, WR achieved the best and worst scores, respectively. Therefore, using only summative metrics, we can conclude that GPT-4 outperforms GPT-3.5 in SpyGame. However, the proposed metrics suggest different results: both GPT-4 and GPT-3.5 are good at deceptive control regarding VR and VE.
@@ V Qualitative Analysis
To reconfirm the findings of quantitative analysis from another perspective, we conducted a thematic analysis. We defined three types of reasoning errors in LLM reasoning while deciding their actions by coding abnormal patterns of 15 randomly selected gameplay logs. As a result, we found that these patterns are related to psychological concepts, including identity confusion, memory distortion, and dissociation[23]. In addition, we found that these concepts are related to the results of the quantitative analysis, especially for caught rate (CR) and guess success (GS). The following subsections illustrate the procedure and results of this qualitative analysis.
@@ V-A Analysis Procedure
$N$* Character ambiguity Memory distortion Dissociation
@@
* The total number of reasoning steps in each condition
@@ V-B1 Exposure
Exposure is a case in which a citizen’s utterance literally reveals the hidden location. The Fleiss’ Kappa value for this category was 0.7875. Note that this pattern can only be identified in utterances rather than in reasoning steps. Therefore, we report the frequency of this category based on utterances; among 729 citizen utterances, 125 were labeled as exposure. The following quote provides a representative example:
@@
We suspect that LLMs may not have sufficient common sense to understand the game rules, causing this pattern. In particular, SpyGame has a rule “citizens should not expose the location to the spy.” A human citizen player can notice that a direct mention to other players of the hidden location during a game causes their defeat. However, some LLMs may not have this common sense among human players; they simply go through the mentions. Therefore, LLMs believe that they can reveal their information to other citizens because the game rule does not algorithmically prohibit such behaviors and the citizens already know that fact. Thus, citizen LLMs sometimes publicly reveal their hidden information.
V-B2 Dissociation
[Player4, the spy, tries to guess the hidden location.]
We suspect that this category occurs because LLMs are not good at differentiating between numerical symbols. As previous studies have reported that LLMs suffer from biases induced by proper names[25], we named each player as ‘Player#’ (e.g., Player1) to avoid such biases. Therefore, when an LLM attempts to collect clues from the conversation log, it has to distinguish the source of each utterance using a single difference between player names: a single digit. However, previous studies have reported that LLMs struggle to recognize different digit tokens during their computation processes [26]. Thus, LLM players can misinterpret their words as those of others in the log, which triggers dissociation.
V-B3 Memory distortion
V-B4 Character ambiguity
Character ambiguity is a case in which the spy confuses their own identity with that of a citizen. This category is named after a similar psychological concept [28]. In an SDG, understanding a player’s identity is related to two major concepts affecting gameplay: (1) the team of the player and (2) the goal of the player. Therefore, we divided abnormal patterns of character ambiguity into two sub-categories: team misunderstanding and goal misunderstanding. These two subcategories can occur independently. We present details of the subcategories in the below paragraphs. In summary, Fleiss’ Kappa for this category was 0.563. Among 506 reasoning steps in the data, we identified 21 steps with character ambiguity.
@@ V-B4 Character ambiguity
We suspect that this category occurs is due to the memory limitation of LLMs. To understand a player’s identity correctly, an LLM should retain information about the player’s identity during the decoding process. Since we provided the game rule, identity of a player, and conversation logs together as an input prompt, the LLM needs to distinguish information about the identity from the input. However, such a distinction is difficult for the LLM because the other parts of the input prompt are much larger than the identity part. For example, the explanation of the game rules takes more than 130 words, whereas that of identity takes only 10 words. Thus, LLM players can confuse their identities with other characters written in the game rules, triggering character ambiguity.
@@ V-C Reconfirmation of Quantitative Analysis
Combining the results of the two analyses confirms the findings of the quantitative analysis. Specifically, we discovered that these quantitative labels, except exposure, can be connected to two abilities in the quantitative analysis: subtext inference and deceptive control.
First, memory distortion and dissociation may hinder a spy from inferring subtext during the game. When we compared the frequency of memory distortion and dissociation with the results of the four subtext inference metrics, the trend of frequency was similar to that of the metrics. For example, against strong citizens, GPT-3.5 showed the highest errors in both memory distortion and dissociation and achieved the lowest performance in subtext inference. Similarly, GPT-4 showed fewer errors while achieving a higher performance than the other models. We suspect that this similarity between trends is due to the quality of information that the spy considers during their reasoning process. The spy player should select and combine pieces of information from the conversation to identify the hidden location. When memory distortion occurs, the spy can regard misinterpreted pieces as appropriate. Similarly, when dissociation occurs, the spy regards apocryphal information that they previously said as information provided by citizens. Therefore, such misunderstood pieces can be obstacles to reaching the correct answer.
Second, character ambiguity increases the likelihood of a spy’s failure of deceptive control. When we compared the frequency of character ambiguity and the results of the three metrics regarding deceptive control, the trend of frequency was similar to that of the metrics. For example, against strong citizens, GPT-3.5 showed the lowest errors in character ambiguity and achieved the highest performance in such metrics. Also, Llama showed more errors while achieving a lower performance than the other models. We suspect that this similarity between the trends is due to the loss of the spy’s need for deception. When character ambiguity occurs in the spy’s reasoning steps, the spy confuses their character with that of a citizen. Such confusion leads the spy to make an unusual utterance, which draws suspicion.
@@ VI Conclusion
To assess whether LLMs can effectively support obscured communication, we addressed two key limitations in prior evaluation methodologies. The first is the reliance on coarse-grained quantitative metrics, which fail to capture fine-grained communicative behaviors. The second is the lack of structured qualitative analyses that can contextualize evaluation results or uncover the reasoning processes underlying model behavior. To investigate these limitations in a controlled setting, we conducted experiments using SpyGame, a variant of Spyfall, and collected gameplay logs from four LLMs. We analyzed these logs through both quantitative and qualitative methods. First, we proposed eight fine-grained metrics designed to evaluate two core abilities: subtext inference and deceptive control. These metrics demonstrated more effective in identifying subtle behavioral differences between models. Second, we performed a thematic analysis to identify reasoning failures not captured by quantitative metrics. This analysis revealed four major categories of failure that align with the patterns observed in the quantitative evaluation, offering complementary insights. Our findings highlight the importance of microscopic, multi-perspective evaluation for understanding how LLMs perform in obscured communicative settings. We hope this study contributes to a deeper understanding on LLMs in complex social interactions.